### v1

# Solvation Free Energies from Neural Thermodynamic Integration

Abstract
We propose to compute solvation free energies via thermodynamic integration along a neural-network potential interpolating between two target Hamiltonians. We use a stochastic interpolant to define an interpolation between the distributions at the level of samples and optimize a neural network potential to match the corresponding equilibrium potential at every intermediate time-step. Once the alignment between the interpolating samples and the interpolating potentials is sufficiently accurate, the free-energy difference between the two Hamiltonians can be estimated using (neural) thermodynamic integration. We validate our method to compute solvation free energies on several benchmark systems: a Lennard-Jones particle in a Lennard-Jones fluid, as well as the insertion of both water and methane solutes in a water solvent at atomistic resolution.
1 Introduction
Estimating free energy differences is at the heart of understanding a wide range of physical, chemical, and biological processes, from protein folding and ligand binding to phase transitions in materials. These calculations offer invaluable insights into the stability of molecular conformations, the spontaneity of chemical reactions, and the mechanisms that drive phase changes (Gao et al., 2006; Mobley & Gilson, 2017; Agarwal et al., 2021). The traditional approaches for performing such calculations, such as thermodynamic integration (TI) and free energy perturbation (FEP) have recently been complemented with ideas from the machine learning community (Noé et al., 2019; Wirnsberger et al., 2020; Invernizzi et al., 2022).
This work generalizes NeuralTI (Máté et al., 2024), a machine-learning enhanced variant of TI, to estimate the free energy difference between two arbitrary Hamiltonians. In §3 we argue that stochastic interpolants provide a natural framework for this setting and describe our approach. Finally, in §4 we validate our methodology by computing solvation free energies on several systems: ($i$) a Lennard-Jones (LJ) solute in an LJ solvent, and the insertion of ($ii$) a water solute and ($iii$) a methane solute in a box of water, both at atomistic resolution. All free energies show excellent agreement with reference calculations.
Figure 1: An interpolating family of distributions coupling a solute (brown) to the solvent (blue). The potentials $U_{A},U_{B}$ and $U_{AB}$ denote the interactions within the solvent, within the solute and between the two components, respectively. The interpolation $U_{t}=U_{A}+U_{B}+tU_{AB}$ is often used in TI calculations to compute the free energy of the coupling of the solute, we include an additional trainable potential $t(1-t)U_{t}^{\theta}$ and train it to be the equilibrium potential of a stochastic interpolant.
2 Background
We begin with fixing some notation. Depending on which arguments are important in the context, we will use $U_{t}(x)$ or $U(t,x)$ or simply $U$ to denote time-dependent energy/potential functions and $\rho_{t}$ for the corresponding Boltzmann probability density $\frac{1}{Z_{t}}e^{-\beta U_{t}}$, with inverse temperature $\beta=1/k_{B}T$. The symbol $\nabla U$ is used for the spatial gradient. Isotropic Gaussian distributions with variance $\sigma^{2}$ are written as $\mathcal{G}_{\sigma}$ and the convolution of two functions, $f$ and $g$ as $f\star g$.
Boltzmann densities and free energies
[…]
Thermodynamic Integration (TI)
Thermodynamic integration (Kirkwood, 1935) is a method for estimating the free energy difference between two Hamiltonians relying on a family of potentials, $U_{\lambda}$, interpolating between $U_{0}$ and $U_{1}$.
$\beta\Delta F_{0\rightarrow 1}$ $=\log Z_{0}-\log Z_{1}=-\int_{0}^{1}\text{d}\lambda\,\partial_{\lambda}\log Z_{\lambda}=-\int_{0}^{1}\text{d}\lambda\,\frac{1}{Z_{\lambda}}\partial_{\lambda}\left(\int\text{d}x\,\text{e}^{-\beta U_{\lambda}(x)}\right)$ (1)
$=\beta\int_{0}^{1}\text{d}\lambda\,\frac{1}{Z_{\lambda}}\left(\int\text{d}x\,\text{e}^{-\beta U_{\lambda}(x)}\partial_{\lambda}U_{\lambda}(x)\right)=\beta\int_{0}^{1}\text{d}\lambda\,\left\langle\partial_{\lambda}U_{\lambda}\right\rangle_{\lambda},$ (2)
where $\langle\partial_{\lambda}U_{\lambda}\rangle_{\lambda}$ denotes the expected value of $\partial_{\lambda}U_{\lambda}$ under $\rho_{\lambda}=Z_{\lambda}^{-1}e^{-\beta U_{\lambda}}$. To numerically estimate the value of the last integral one needs an interpolating potential function $U_{\lambda}$ and samples from the corresponding interpolating equilibrium distributions $\rho_{\lambda}$. Usually one first fixes an interpolation $U_{\lambda}$ and performs Monte Carlo sampling or molecular dynamics (MD) simulations at intermediate $\lambda$ values to end up with the estimate
$\Delta\hat{F}_{0\rightarrow 1}=\frac{1}{N+1}\sum_{\lambda_{i}}\,\mathbb{E}_{x\sim\rho_{\lambda_{i}}}\big{[}\partial_{\lambda}U_{\lambda}(x)\big{]}\qquad\lambda_{i}\in\{0,1/N,2/N,...,1\}$ (3)
Neural TI
Máté et al. (2024) proposed to parametrize a time-dependent energy function $U_{t}$ and use denoising score matching to align the force $\nabla U_{t}$ with the scores of increasingly noised versions of the target Boltzmann distribution $e^{-\beta U_{0}}/Z_{0}$. As the noise level $\sigma(t)$ starts small at $t=0$, and is large at $t=1$, the trained $U_{t}$ provides and interpolating potential between the target and a tractable prior. Since samples from the corresponding Boltzmann distributions $e^{-\beta U_{t}}\approx\mathcal{G}_{\sigma_{t}}\star e^{-\beta U_{0}}$ are also (approximately) available, TI can be performed along $U_{t}$ to estimate the free-energy difference between a trivial and a non-trivial distribution. In this case, these were matched to the ideal gas and the Boltzmann distribution of the fully coupled LJ liquids, so as to associate the resulting free energy to the coupling of all the interactions.
Traditional approaches to solvation free-energy calculations often consist of first fixing the interpolation in the space of potentials (e.g., linear coupling), to subsequently generate samples from Monte Carlo sampling or molecular dynamics simulations Mey et al. (2020). Here instead, we first fix the sampling procedure and learn the interpolating family of potentials.
3 Method
The goal of this work is to extend the neural TI framework to the estimation of free energy differences between two non-trivial Hamiltonians. To this end, we consider a pair of potential functions, $U_{0},U_{1}:\mathcal{M}\rightarrow\mathbb{R}$ on a manifold $\mathcal{M}$ representing the configuration space of some system. In our experiments, we model particle systems and rigid molecules in a periodic box. For the particle system $\mathcal{M}$ will be the $3N$-torus, $\mathbb{T}^{3N}$ representing the positions of $N$ particles in a 3-torus. For the rigid molecular system, the configuration of each molecule is given by its position in $\mathbb{T}^{3}$ and its orientation, thus the configuration space of the full system is $\mathcal{M}=[\mathbb{T}^{3}\times SO(3)]^{N}$.
Stochastic Interpolants
We assume that we can sample $x_{0}\sim\rho_{0}\propto e^{-\beta U_{0}}$ and $x_{1}\sim\rho_{1}\propto e^{-\beta U_{1}}$. We then define an interpolating family of distributions $\rho_{t}$ for $10<t<$ 1 by constructing a procedure to sample $x_{t}$ using a stochastic interpolant (Albergo & Vanden-Eijnden, 2023; Albergo et al., 2023). For the sake of simplicity, let us first assume that $\mathcal{M}=\mathbb{R}^{n}$. A stochastic interpolant generates interpolating samples from $\rho_{t}$ by
$x_{t}=I(x_{0},x_{1},t)+\sigma_{t}z$ (4)
where $x_{0}\sim\rho_{0},x_{1}\sim\rho_{1},z\sim\mathcal{N}(0,1)$, $I$ satisfies $I(x_{0},x_{1},0)=x_{0},I(x_{0},x_{1},1)=x_{1}$, e.g., $I(x_{0},x_{1},t)=(1-t)x_{0}+tx_{1}$ and $\sigma_{0},\sigma_{1}\ll 1$. If $\sigma_{0}=\sigma_{1}=0$, then the boundary conditions are exactly satisfied, but $0<\sigma_{0},\sigma_{1}\ll 1$ also allows score-based training, while only slightly violating the boundary conditions. Conditioned on $x_{0}$ and $x_{1}$, $\rho_{t}(x_{t}|x_{0},x_{1})$ is an isotropic Gaussian with mean $I(x_{0},x_{t},t)$, variance $\sigma_{t}$ and Stein-score $\nabla\log\rho_{t}(x_{t}|x_{0},x_{1})=\frac{1}{\sigma_{t}}(I(x_{0},x_{1},t)-x_{t})=-z/\sigma_{t}$. The marginal score $\nabla\log\rho_{t}(x)$ reads
$\nabla\log\rho_{t}(x)=\mathbb{E}_{x_{0},x_{1}}\left[\nabla\log\rho_{t}(x_{t}|x_{0},x_{1})\right]=\mathbb{E}_{x_{0},x_{1}}\left[-z/\sigma_{t}\right],$ (5)
enabling the training of an energy-based model using the usual denoising score matching (DSM) objective (Vincent, 2011). This approach can be naturally extended to non-Euclidean $\mathcal{M}$ by working in local charts on the manifold (De Bortoli et al., 2022).
Rolling estimate of $F_{0\to 1}$
The DSM objective requires the computation of the spatial gradient $\nabla_{x}U_{t}(x)$ at samples $x_{t}$, sampled at $(t,x)\sim(\mathcal{U}([0,1]),\rho_{t})$. The temporal derivative of the energy function, $\partial_{t}U_{t}(x)$, can be easily calculated thanks to automatic differentiation at the same points, yielding a Monte Carlo estimate of the TI integral in Equation (2),
$\Delta\hat{F}_{0\rightarrow 1}=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x\sim\rho_{t}}\big{[}\partial_{t}U_{t}(x)\big{]}$ (6)
In our experiments, at every training step we compute both the spatial and temporal derivatives of $U_{t}$ and use $\nabla U_{t}$ for the DSM objective and $\partial_{t}U_{t}$ for a rolling estimate of the free-energy difference.
Locally consistent potentials
Our strategy relies on learning a time-dependent potential $U_{t}$ corresponding to the samples $x_{t}\sim\rho_{t}$ produced by a stochastic interpolant. To perform TI, the samples $x_{t}$ and the energy function $\beta U_{t}$ need to be compatible with each other in the sense that $x_{t}\sim e^{-\beta U_{t}}$. An energy-based model trained with the DSM objective can only fit an energy function up to an additive constant. Since we aim to learn a time-dependent family of energies, the degrees of freedom correspond to a time-dependent function defining the energy scale along the trajectory. Because we enforce the boundary conditions by construction (equation 12), this degree of freedom vanishes when integrated over the full time interval $[0,1]$.
A possible issue arises for a target domain that is disconnected or a distribution that is multimodal: the additive constant becomes local. Thus, strictly speaking, our model does not learn a function $U_{t}$ such that $e^{-\beta U_{t}}$ is globally proportional to $\rho_{t}$, but only locally proportional. In other words, it does not preserve the relative weights of the modes, but does preserve the relative likelihoods within the modes. (Song & Ermon, 2019) solve this by learning a family of models with increasing noise levels, making use of the fact that for high perturbations the target distribution becomes connected. Alternatively, Gutmann & Hyvärinen (2010) overcome the issue by contrasting the data with a second distribution of known density. In the present work, the enforced boundary conditions alleviate the issue: All offsets introduced to the relative weighting of the modes also integrate to zero over the time interval $[0,1]$. See Figure 2 for a one-dimensional illustration of this argument.
Figure 2: Unnormalized interpolating densities $e^{-\beta U_{t}^{\theta}}$ learned from different initializations. The two target potentials are given by a standard Gaussian at $t=0$ ($\log Z_{0}=1$) and the sum of two Gaussian densities at $t=1$ ($\log Z_{1}=2$). The 4 different models learn 4 different relative weightings of the modes along the interpolation, but have a consistent prediction for the free energy difference of $\log 2\approx 0.69$.
4 Experiments
We evaluate the proposed method by estimating solvation free energies, i.e., we concentrate on the case where the potentials $U_{0},U_{1}$ are of the form
$U_{0}$ $=U_{\mathrm{solvent}}+U_{\mathrm{solute}}$ (7)
$U_{1}$ $=U_{\mathrm{solvent}}+U_{\mathrm{solute}}+U_{\mathrm{solute-solvent}}$ (8)
where $U_{\mathrm{solvent}},U_{\mathrm{solute}}$ and $U_{\mathrm{solute-solvent}}$ represent the interactions within the solvent, within the solute and cross interactions between the two components, respectively.
4.1 Solute solvation of a Lennard-Jones fluid
System description
We consider a system of 512 solvent particles and an additional solute particle with Lennard-Jones (LJ) interactions between them. The LJ parameters of the solvent are denoted by $(\varepsilon_{A},\sigma_{A})$ while for the solute $(\varepsilon_{B},\sigma_{B})=(2\varepsilon_{A},2\sigma_{A})$ and the LJ interactions are computed by the Lorentz-Berthelot rules (Lorentz, 1881; Berthelot, 1898) $(\varepsilon_{AB}=\sqrt{\varepsilon_{A}\varepsilon_{B}},\sigma_{AB}=(\sigma_{A}+\sigma_{B})/2)$. The simulation box is of size $9\sigma_{A}\times 9\sigma_{A}\times 9\sigma_{A}$ with periodic boundary conditions. The potential functions $U_{\mathrm{solvent}},U_{\mathrm{solute}}+U_{\mathrm{solute-solvent}}$ are then given by
$U_{\mathrm{solute}}$ $=0$ (9)
$U_{\mathrm{solvent}}$ $=\sum_{i\neq j}4\varepsilon_{A}\left[\left(\frac{\sigma_{A}}{r_{ij}}\right)^{12}-\left(\frac{\sigma_{A}}{r_{ij}}\right)^{6}\right]$ (10)
$U_{\mathrm{solute-solvent}}$ $=\sum_{i}4\varepsilon_{AB}\left[\left(\frac{\sigma_{AB}}{r_{i}}\right)^{12}-\left(\frac{\sigma_{AB}}{r_{i}}\right)^{6}\right]$ (11)
where $r_{ij}$ is the distance between the solvent particles with indices $i$ and $j$, and $r_{i}$ is the distance between the solute and the solvent particle with index $i$.
Ansatz for $U_{t}$
We parameterize the interpolating potential in a way that enforces the boundary conditions at $t\in\{0,1\}$ (Máté & Fleuret, 2023).
$U_{t}(x)=b_{t}^{A}U_{\mathrm{solvent}}(x,a_{t}^{A})+b_{t}^{B}U_{\mathrm{solute}}(x,a_{t}^{B})+b^{AB}_{t}U_{\mathrm{solvent-solute}}(x,a^{AB}_{t})+b^{\mathrm{U}}_{t}U_{t}^{\theta}(x)$ (12)
where $U_{t}^{\theta}(x)$ is the same graph neural network as in Máté et al. (2024) with node features encoding the particle type (solute or solvent) of the nodes. The functions $a^{A}_{t},a^{B}_{t},a^{AB}_{t}$ denote softening parameters to respective components of the potential (Beutler et al., 1994). We use the following softened form of the LJ-potential
$U(r,a)=4\varepsilon\left[\left(\frac{\sigma^{2}}{a\sigma^{2}+r^{2}}\right)^{6}-\left(\frac{\sigma^{2}}{a\sigma^{2}+r^{2}}\right)^{3}\right],$ (13)
which corresponds to evaluating the Lennard-Jones potential at $\sqrt{r^{2}+a\sigma^{2}}$. When $a=0$, no softening is applied, and larger values of $a$ mean larger softening deformation of the original Lennard-Jones interactions. We enforce the boundary conditions of the the $b_{t}$ and $a_{t}$ functions by learning them with parametrization
$b_{t}^{A}=[1-t(1-t)]e^{\beta_{t}^{\mathrm{A}}}\quad b_{t}^{B}=[1-t(1-t)]e^{\beta_{t}^{\mathrm{B}}}\quad b_{t}^{\mathrm{AB}}=te^{t\beta_{t}^{\mathrm{AB}}}\quad b_{t}^{\mathrm{U}}=t(1-t)e^{\beta_{t}^{\mathrm{U}}}$ (14)
$a_{t}^{\mathrm{A}}=t(1-t)e^{\alpha_{t}^{\mathrm{A}}}\quad a_{t}^{\mathrm{B}}=t(1-t)e^{\alpha_{t}^{\mathrm{B}}}\quad a_{t}^{\mathrm{AB}}=(1-t)e^{t\alpha_{t}^{\mathrm{AB}}}$ (15)
where the $t\mapsto\alpha_{t},t\mapsto\beta_{t}$ functions are learned by a small neural network.
Regularization of $\partial_{t}U_{t}$
The function we are aiming to learn is $U_{t}(x)$. The DSM objective controls its behavior with respect to variations of $x$, but there is no training signal controlling the behaviour of $U_{t}(x)$ with respect to changes to the temporal argument. To avoid large absolute values of $\partial_{t}U_{t}(x)$ we also include the $L_{2}$-norm of $\partial_{t}U_{t}$ to our optimization objective. Our training objective is thus
$\mathcal{L}(\theta)$ $=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x_{t}\sim\rho_{t}}\Big{[}\mathcal{L}_{\mathrm{DSM}}+\kappa\mathcal{L}_{\mathrm{reg}}\Big{]}$ (16)
$\mathcal{L}_{\mathrm{DSM}}=\sigma_{t}^{2}||\nabla_{x}U_{t}(x_{t})-z||^{2}$ (17)
$\mathcal{L}_{\mathrm{reg}}=\frac{1}{2}(\partial_{t}U_{t}(x_{t}))^{2}$ (18)
where $U_{t}$ depends on $\theta$ through Equation 12. Since $\partial_{t}U_{t}$ is the very thing we are interested in measuring, we use a small relative weight for this regularization in our experiments, $\kappa=10^{-10}$. The goal is not minimize the $L_{2}$-norm of $\partial_{t}U_{t}$, but to prevent it from diverging.
Interpolation
Given samples $x_{0},x_{1}\sim(\rho_{0},\rho_{1})$, we construct the function $I(x_{0},x_{1},t)$ in Equation (4) by translating both $x_{0}$ and $x_{1}$ to center the solute and then compute the optimal transport (OT) between the solvent particles with respect to the squared toroidal distance between the solvent particles. Note that we compute the OT-pairings between the solvent particles from two given snapshots $x_{0},x_{1}$ and not between datapoints in a batch as proposed by Tong et al. (2023) During the interpolation, the solvent particles move along the toroidal geodesics given by the OT-pairing , and we keep the solute frozen.
Figure 3: Solute solvation of a Lennard-Jones fluid. Left: Solvation free energy estimates by TI and neural TI as a function of number of reference simulations and training steps, respectively. Center: Detailed view at later training stages. Softening parameters of the LJ-interactions after training (right).
Results
We compare our estimates with TI using evenly spaced time-slices along the time-interval $[0,1]$. We perform 51 simulations in total at $t\in\{0.00,0.02,0.04,...,1.00\}$, and find that the solvation free energy lies between $-21k_{B}T$ and $-19k_{B}T$. The rolling estimate of neural TI converges to $\approx-19.5k_{B}T$. The time evolution of these values are shown in the left subplot of Figure 3. The right subplot of Figure 3 displays the learned softening schedule of the solvent-solvent and solute-solvent Lennard-Jones interactions. The learned softening functions show that our interpolation does not correspond to the usual linear interpolation $(1-t)U_{0}+tU_{1}$, even the interactions within the solvent are softened and decoupled along the trajectory.
4.2 Hydration free energies
To move to more realistic chemical systems, we now extend the experimental setup to compute hydration free energies of small molecules, namely water and methane, from atomistic simulations. We work with rigid molecules, i.e., we consider Lennard-Jones and Coulomb, but not bonded, interactions. Microstates of single molecules are thus described by the the position of their central atom (oxygen for water, carbon for methane) and their orientation, i.e., a group element of $SO(3)$.
System description
The solvent consists of 216 TIP4P water molecules (Jorgensen et al., 1983) at density 1 g/cm3, resulting in a simulation box of size $18.62{\AA}\times 18.62{\AA}\times 18.62{\AA}$. The hydration of water—solvation of water in water—simply consists of coupling one extra TIP4P molecule in the box. For methane, we rely on the atomistic OPLS (OPLS-AA) force field (MacKerell Jr et al., 1998). Coulomb interactions are modelled with the reaction field method (Tironi et al., 1995)
$E_{\text{Coulomb}}(r)$ =$=\begin{cases}k_{C}{q_{1}q_{2}}\left(\frac{1}{r}+k_{\text{rf}}r^{2}-c_{\text{rf}}\right),\quad\text{if}\quad r\leq r_{\text{cutoff}}\\ 0,\quad\text{if}\quad r>r_{\text{cutoff}}\end{cases}$ (19)
$k_{\text{rf}}$ $=\left(\frac{1}{r_{\text{cutoff}}^{3}}\right)\left(\frac{\epsilon_{\text{water}}-1}{2\epsilon_{\text{water}}+1}\right)$ (20)
$c_{\text{rf}}$ $=\left(\frac{1}{r_{\text{cutoff}}}\right)\left(\frac{3\epsilon_{\text{water}}}{2\epsilon_{\text{water}}+1}\right)$ (21)
where $r_{\text{cutoff}}$ is the cutoff radius, which we set to $8.5\AA$, $k_{C}$ is Coulomb’s constant, $q_{1},q_{2}$ are the charges of the atoms, $r$ is the distance between them and $\epsilon_{\text{water}}$ is the dielectric constant of water, $\epsilon_{\text{water}}=78$. To soften this potential with softening parameter $a$, we define
$E^{\text{soft}}_{\text{Coulomb}}(r,a)=E_{\text{Coulomb}}(\sqrt{r^{2}+a})$ (22)
To learn the interpolating potential we use the ansatz of equation (12), which now contains both the soft Coulomb interactions defined by equation (22) in addition to the soft Lennard-Jones interactions defined by equation (13). Importantly, unlike the usual practice of TI, we do not prescribe to couple the Lennard-Jones interactions first, and the Coulombic interactions later, we let the network learn the complete coupling schedule with separate softening parameters for the two interactions.
Interpolation
Each molecule is represented by the position of its central atom (oxygen for water, carbon for methane) in $\mathbb{T}^{3}$ and its orientation in $SO(3)$. To interpolate between two snapshots, $x_{0}$ and $x_{1}$, we perform the same steps as in the previous section for the central atoms, and use the geodesic interpolation on $SO(3)$ between the orientations. To further minimize the rotational transport cost, we use the fact that both the rigid water and methane molecules have non-trivial symmetric groups, $\mathbb{Z}_{2}$ and $S_{4}$, respectively. Applying any of these transformation does not change the physical system itself, only its numerical representation [^1] Geometrically speaking, this means that the configuration space of a single rigid molecule can be reduced from $\mathbb{T}^{3}\times SO(3)$ to $\mathbb{T}^{3}\times SO(3)/G$, where $G\in\{\mathbb{Z}_{2},S4\}$ is the inner symmetry group of the molecule. thus we compute the geodesic distances corresponding to the physically equivalent orientations and choose the one with the minimal geodesic distance.
Architecture
[…]
Results
We benchmark our calculations on previously reported calculations of the hydration free energy for both water and methane: $-10.7k_{B}T$ (Jorgensen et al., 1989) and $3.4k_{B}T$ (Mobley et al., 2009), respectively. Figure 4 (right panel) compares our rolling neural TI estimates (equation (6)) with the reference values mentioned. For each system we provide two individual calculations with different initial random seeds. For the moving average we use a window size of 4,000 batches of batch size 8, i.e., 32,000 samples. The runs for water and methane use the same architecture of the neural network potential, same configuration, the runs only differ by the solute. The runs for the water and methane visibly diverge after 5,000 traing steps and stay in the ±$\pm 4k_{B}T$ range of the experimental value after 10,000 training steps. The estimates for water appear to be more consistent, this is likely due to the fact that the water-in-water system is more homogeneous than the methane-in-water system. All runs were performed on an Nvidia 3090 and took about 4 hours to reach the plotted 50,000 training steps.
Figure 4: Illustration of the coupled water (left) and methane (center) molecules. The oxygen of the solute water molecule is colored blue to distinguish it from the solvent. Hydration free energy estimates of water and methane in TIP4P water (right). The dashed lines denote the experimental hydration free energies and the the curves of the same our rolling estimate from 2 different random seeds for each solute.
5 Conclusion
We build on the ideas introduced by Máté et al. (2024) to estimate free energy differences between arbitrary pairs of target distributions. This means that diffusion models are replaced by their generalization, stochastic interpolants (Albergo & Vanden-Eijnden, 2023), a natural framework for our purposes. The method now relies on two, possibly non-trivial, Hamiltonians over the same configuration space, samples from their Boltzmann distributions, and an interpolating procedure between samples from these two distributions. The upside of this approach, compared to traditional TI, is that we only need data at the end-point Boltzmann distributions, entirely alleviating the need for intermediate reference MC or MD simulations.
Additionally, we streamlined the process by incorporating techniques like the rolling estimate of free energy differences (equation 6) and the regularization of the temporal derivative (equation 16), both of which we found beneficial during our experiments. In §3 we provide some theoretical insight into the characteristics of the learning paradigm.
The method was verified by accurately estimating the solvation free energy of a Lennard-Jones solute in a Lennard-Jones solvent and the hydration free energies of water and methane. We restricted our attention to rigid molecules, future work could explore how to include bonded interactions into the approach.
Acknowledgments
We thank Aleksander Durumeric, Daniel Nagel and Youssef Saied for discussions. BM acknowledges financial support by the Swiss National Science Foundation under grant number CR - SII5 - 193716 (RODEM). TB acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).